# Rajamanickam, Almutari and Majumdar

### v2

@@ Abstract
We study the effects of colloidal nanoparticles (NPs) in cholesteric liquid crystal samples in the dilute limit, in a Landau–de Gennes theoretical framework. The effects of the suspended NPs are captured by a homogenized energy, as outlined in [7]. For spatially homogeneous samples, we explicitly compute the critical points and minimizers of the modified Landau–de Gennes energy and show that the presence of NP eliminates the first-order isotropic-nematic phase transition, stabilises elusive biaxial phases over some temperature ranges, and that the symmetry of the NP boundary conditions or surface treatments dictates the bulk equilibrium phase at high temperatures. We also numerically demonstrate structural transitions from twisted helical director profiles to untwisted director profiles in cholesteric-filled channel geometries, driven by the collective effects of the NPs and increasing temperature. These transitions are reversible upon lowering the temperature in sufficiently large domains, where thermal hysteresis can also be observed. This behaviour opens interesting avenues for tuning the optical properties of confined, nano-doped cholesteric systems.
@@
Landau–de Gennes theory; homogenization theory; nanoparticles; cholesterics; biaxiality.
@@ 1 Introduction
There are three canonical LC phases: the simplest nematic (NLC) phase with no positional ordering but with long-range orientational ordering characterised by nematic directors; the smectic phase with orientational and partial positional ordering, i.e., the smectic molecules arrange themselves in layers that are stacked on top of each other, and the cholesteric (CLC) phase, which is similar to the NLC phase except that the cholesteric director naturally twists in space, rendering stable helical director profiles with exceptional optical properties [11]. NLCs are the working material of choice for a variety of electro-optic devices, e.g., the multi-billion dollar liquid crystal display industry, largely because of their strong responses to light and electric fields. CLCs exhibit distinctive helical director profiles and their natural twistedness, or chirality, has found many uses in photonics, sensors, anti-counterfeiting and bio-inspired systems  [17, 23]. There is substantial contemporary interest in doping NLC and CLC systems with colloidal nanoparticles (NPs) to tune their optical and mechanical properties by means of NP shapes, sizes, distributions and surface treatments on NP-LC interfaces [16, 22, 23, 15, 31]. For example, in [25], the authors experimentally and theoretically investigate a dilute suspension of quantum dots in a bent-core nematic system, presenting detailed experimental results on optical textures, phase transition temperatures and dielectric anisotropy data of these doped systems. These quantum dots, typically synthesised using Cd and Zn-based compounds, are capped with suitable ligands to prevent aggregation, making them good candidates for suspension in the LC matrix [25]. The authors demonstrate that the quantum dots and, NPs in general, can be used to tune the dielectric anisotropy, phase transition temperatures and optical textures of confined bent-core nematic systems [25]; related works can also be found in [2, 22, 21, 27, 16] with regards to the effects of NPs on operating voltages, elastic constants and response times in LC systems. In other words, doping offers exciting prospects for tailor-made or customised LC systems on demand, as well as tunable metamaterials in general.
@@
In this paper, we study a prototype system comprising a dilute suspension of colloidal spherical NPs in a CLC sample, although many of our conclusions will naturally extend to NLC systems. In a dilute suspension, the NP size is much smaller than the average distance between NPs, which in turn is much smaller than the macroscopic sample length scales. The total volume fraction of the NPs is much less than unity, and in this dilute limit, the collective effects of the NPs can be effectively homogenized into a bulk effect. We work in the celebrated Landau–de Gennes (LdG) theoretical framework for NLCs and CLCs, wherein the macroscopic NLC or CLC state is described by the macroscopic $\mathbf{Q}$-tensor order parameter [11, 24, 18]. The $\mathbf{Q}$-tensor order parameter is relatable to experimentally measurable quantities such as dielectric anisotropy and magnetic susceptibility, which characterize the LC’s response to external electric and magnetic fields [11]. Mathematically speaking, the LdG $\mathbf{Q}$-tensor is a symmetric, traceless $3\times 3$ matrix, whose eigenvectors model the LC directors, i.e., the special directions of averaged molecular alignment, and whose corresponding eigenvalues are scalar order parameters that quantify the degree of orientational ordering. A LC phase is uniaxial if it has a single distinguished material direction and is biaxial if it has two special directions; more details are given in the next section. The LdG theory has an associated free energy, typically of the form
@@
in the absence of external fields or surface effects. Physically observable configurations are modelled by the LdG energy minimizers subject to the problem constraints. In the case of spatially homogeneous samples, the LdG energy reduces to the bulk potential, $f_{B}$, which is typically a polynomial in the scalar invariants of $\mathbf{Q}$ [18]. The bulk equilibrium phases are then modelled in terms of the minimizers of $f_{B}$, which can be computed explicitly as a function of temperature [24]. The bulk potential, $f_{B}$, is identical for NLC and CLC phases. The elastic energy density, $f_{el}$, depends on both $\mathbf{Q}$ and its gradients, and penalises spatial inhomogeneities. In fact, $f_{el}$, has different forms for the NLC and CLC phases because the natural ground state for a CLC is a helical director profile, whereas that for a NLC is a uniform, constant director profile. The critical points and the minimizers of $\pazocal{F}$ are classical solutions to the associated Euler-Lagrange equations: a system of nonlinear, coupled partial differential equations, subject to prescribed constraints.
The question of interest is: how is the LdG energy modified by the NPs in the dilute limit described above? This question has been answered in [7, 8]. In the dilute limit, the NPs do not interact with each other and the NP-LC interaction is captured by a surface energy, $f_{s}$, that depends on the NP geometry and the surface treatment on the NP-LC interface, $\Sigma$. The collective effects of the NPs is, informally speaking, the sum of the surface energies over all NP-LC interfaces, which yields a homogenized bulk energy, $f_{hom}(\mathbf{Q})$. More precisely, the total LdG energy, including the effects of the suspended NPs is given by
$\pazocal{F}_{\epsilon}[\mathbf{Q}]:=\pazocal{F}[\mathbf{Q}]+\epsilon^{\gamma}\int_{\Sigma}f_{s}\left(\mathbf{Q},\mathbf{Q}_{\nu},\boldsymbol{\nu}\right)~dS$
where $\epsilon$ is related to the NP size, $\gamma$ is an appropriate positive scaling factor determined by the dilute limit, $\nu$ is the normal to the NP-LC interface and $\mathbf{Q}_{\nu}$ is a preferred boundary condition on $\Sigma$. In [7], the authors show that minimizers of $\pazocal{F}_{\epsilon}$ converge to minimizers of
@@
We apply this homogenization framework to a dilute suspension of colloidal NPs in a CLC sample and analyse the properties of the NP-doped CLC systems in terms of the minimizers of $\pazocal{F}_{0}$, as introduced above. In Section 2, we compute the bulk equilibrium phases of these NP-doped systems as functions of the NP properties and temperature. The bulk phases are modelled by the minimizers of the modified bulk potential, which is the sum of $f_{B}$ and $f_{hom}$. In fact, the bulk equilibrium phases do not distinguish between NLC and CLC samples and hence the conclusions in Section 2 apply to both NLC and CLC samples. The NP properties are captured by two parameters: an anchoring coefficient, $W$, on the NP-LC interfaces and a phenomenological matrix, $\mathbf{X}$, that can be computed from $f_{s}$ directly. We find that NPs stabilise elusive biaxial phases in certain temperature ranges, which are typically inaccessible in pure NLC or CLC samples. Equally importantly, the collective symmetry of the NPs (modelled by $\mathbf{X}$) has a direct bearing on the bulk equilibrium phase at high temperatures and breaks the inherent rotational symmetries of the NLC and CLC phases. These concepts are beautifully illustrated by means of phase diagrams in Section 3, which numerically capture how the bulk equilibrium phases depend on $\mathbf{X}$, $W$ and temperature. In Section 4, we apply this framework to a spatially inhomogeneous CLC-filled channel geometry, doped with spherical NPs in the dilute limit. In the absence of NPs, CLCs naturally relax into twisted helical profiles for relatively large channel widths. The NPs induce untwisting transitions in these geometries as temperature increases, i.e., the NPs naturally break the helical symmetry of the CLC phase. In other words, the NPs collectively impose a special direction (determined by $\mathbf{X}$) on the CLC phase, for sufficiently large temperatures and large channel geometries. These temperature-induced untwisting transitions are specific to CLCs. However, using the same arguments, we do expect NPs to induce symmetry-breaking transitions in other confined LC systems too. We conclude with some perspectives in Section 6.
@@ 2 Homogenized bulk energy
We study a dilute system of suspended colloidal NPs, e.g., quantum dots, in a cholesteric liquid crystal-filled channel geometry defined by
@@
with $H\ll L\sim D$. We work in the celebrated phenomenological Landau–de Gennes (LdG) framework [10, 11]. The LdG framework describes the macroscopic state of the cholesteric liquid crystal in terms of the LdG $\mathbf{Q}$-tensor: a symmetric, traceless $3\times 3$ matrix that belongs to the space
@@
We often refer to the eigenvector of $\mathbf{Q}$ with the largest positive eigenvalue as the cholesteric director, i.e., the molecules tend to align along the director or exhibit long-range orientational ordering along the director, in an averaged sense.
A cholesteric phase is said to be isotropic when $\mathbf{Q}=0$ or equivalently, when the eigenvalues of $\mathbf{Q}$ are zero and there are no special directions/directors. In the uniaxial phase, $\mathbf{Q}$ has two degenerate non-zero eigenvalues and can be written in the form
@@
As with most variational theories in materials science, the cholesteric phase has an associated LdG free energy and the equilibrium or physically observable configurations are modelled by LdG energy minimizers, subject to imposed boundary conditions [18]. As mentioned in the Introduction, in the absence of any external fields or surface energies, the LdG energy has two essential contributions:
@@
where $f_{B}$ is the bulk potential, which depends only on the eigenvalues of the $\mathbf{Q}$-tensor and $f_{el}$ is the elastic energy density, which depends on both $\mathbf{Q}$ and its gradient $\nabla\mathbf{Q}$. The bulk potential determines the cholesteric phase (isotropic, uniaxial, or biaxial) in spatially homogeneous samples, as a function of temperature for thermotropic systems. The elastic energy density penalises spatial inhomogeneities stemming from boundary effects, geometric frustration and topological constraints.
@@
where $A=\alpha(T-T^{*})$ is a rescaled temperature, and $B$ and $C$ are positive material-dependent constants. The units of $A,B,C$ are $Nm^{-2}$. Here, $\alpha$ is positive and $T^{*}$ is a critical temperature such that the isotropic phase becomes unstable for $A<0$. As will be shown in the next section, the critical points of $f_{B}$ can be explicitly computed as a function of $T$; in fact, the minimizer of $f_{B}$ models the bulk equilibrium phase at a given temperature for a spatially homogeneous system. The minimizer of $f_{B}$ is the isotropic phase, $\mathbf{Q}=0$, at high temperatures, whereas it is an ordered uniaxial phase for low temperatures. Note that the bulk energy density is identical for nematic and cholesteric liquid crystal phases since cholesterics can be viewed, in non-technical terms, as twisted nematics.
@@
In the homogenization problem, the cholesteric LC molecules are much smaller than the NPs and the NP-LC interaction can be described by a surface energy. The homogenised energy is the collective contribution of the NP-LC surface energies, summed over all the NPs. More precisely, we model the NP-LC interaction energy per unit area by the familiar weak anchoring energy density [24],
@@
In general, we can treat the $\mathbf{X}$-tensor as a phenomenological $3\times 3$, symmetric, traceless tensor, that is purely determined by the NP-LC surface interactions and represents the collective influence of the NPs, even without knowledge of the underlying microscopic details. For example, if $\mathbf{Q}_{\nu}$ corresponds to homeotropic or radial boundary conditions on a spherical particle, then $\mathbf{X}=0$ [7]. For arbitrary $\mathbf{Q}_{\nu}$ and spherical NPs, $\mathbf{X}$ can be non-trivial and we consider multiple possibilities for $\mathbf{X}$ in the next sections. As noted in [7, 8], one can sometimes write $\mathbf{X}=\mathbf{E}\otimes\mathbf{E}$ where $\mathbf{E}$ is an external field and we can redefine $\mathbf{X}=\mathbf{E}\otimes\mathbf{E}-\tfrac{1}{3}\mathbf{I}E^{2}$ to ensure that the isotropic part does not enter into the free energy (8).
@@
where $\pazocal{A}=A/C$, $\pazocal{B}=B/C$ and $W=na^{2}w/C$. Using the values $B=3.7\times 10^{5}$ $Nm^{-2}$ and $C=2.4\times 10^{5}$ $Nm^{-2}$ from  [13] and assuming $w=5\times 10^{-3}Nm^{-1}$, $a=50$ nm and $n=10^{22}$$m^{-3}$, we obtain $\pazocal{B}=1.5$ and $W=0.5$, which will be used in the numerical computations.

### v3

@@ Abstract
We study the effects of colloidal nanoparticles (NPs) in liquid crystal samples in the dilute limit, in a Landau–de Gennes theoretical framework. The effects of the suspended NPs are captured by a homogenized energy, as outlined in [7]. For spatially homogeneous samples, we explicitly compute the critical points and minimizers of the modified Landau–de Gennes energy and show that the presence of NP eliminates the first-order isotropic-nematic phase transition, stabilises elusive biaxial phases over some temperature ranges and that the symmetry of the NP boundary conditions or surface treatments dictates the bulk equilibrium phase at high temperatures. We also numerically demonstrate structural transitions from twisted helical director profiles to untwisted director profiles in cholesteric-filled channel geometries, driven by the collective effects of the NPs and increasing temperature. These transitions are reversible upon lowering the temperature in sufficiently large domains, where thermal hysteresis can also be observed. This behaviour opens interesting avenues for tuning the optical properties of confined, nano-doped cholesteric systems.
@@
Landau–de Gennes theory; homogenization theory; colloidal nanoparticles; cholesterics; biaxiality.
@@ 1 Introduction
There are three canonical LC phases: the simplest nematic (NLC) phase with no positional ordering but with long-range orientational ordering characterised by nematic directors; the smectic phase with orientational and partial positional ordering, i.e., the smectic molecules arrange themselves in layers that are stacked on top of each other and the cholesteric (CLC) phase, which is similar to the NLC phase except that the cholesteric director naturally twists in space, rendering stable helical director profiles with exceptional optical properties [11]. NLCs are the working material of choice for a variety of electro-optic devices, e.g., the multi-billion dollar liquid crystal display industry, largely because of their strong responses to light and electric fields. CLCs exhibit distinctive helical director profiles and their natural twistedness, or chirality, has found many uses in photonics, sensors, anti-counterfeiting and bio-inspired systems  [16, 22]. There is substantial contemporary interest in doping NLC and CLC systems with colloidal nanoparticles (NPs) to tune their optical and mechanical properties by means of NP shapes, sizes, distributions and surface treatments on NP-LC interfaces [15, 21, 22, 14, 31]. For example, in [24], the authors experimentally and theoretically investigate a dilute suspension of quantum dots in a bent-core nematic system, presenting detailed experimental results on optical textures, phase transition temperatures and dielectric anisotropy data of these doped systems. These quantum dots, typically synthesised using Cd and Zn-based compounds, are capped with suitable ligands to prevent aggregation, making them good candidates for suspension in the LC matrix [24]. The authors demonstrate that the quantum dots and, NPs in general, can be used to tune the dielectric anisotropy, phase transition temperatures and optical textures of confined bent-core nematic systems [24]; related works can also be found in [2, 21, 20, 26, 15] with regards to the effects of NPs on operating voltages, elastic constants and response times in LC systems. In other words, doping offers exciting prospects for tailor-made or customised LC systems on demand, as well as tunable metamaterials in general.
@@
In this paper, we study a prototype system comprising a dilute suspension of colloidal spherical NPs in a LC sample (nematics or cholesterics), although many of our conclusions will naturally extend to generic LC systems. In a dilute suspension, the NP size is much smaller than the average distance between NPs, which in turn is much smaller than the macroscopic sample length scales. The total volume fraction of the NPs is much less than unity and in this dilute limit, the collective effects of the NPs can be effectively homogenized into a bulk effect. We work in the celebrated Landau–de Gennes (LdG) theoretical framework for NLCs and CLCs, wherein the macroscopic NLC or CLC state is described by the macroscopic $\mathbf{Q}$-tensor order parameter [11, 23, 17]. The $\mathbf{Q}$-tensor order parameter is relatable to experimentally measurable quantities such as dielectric anisotropy and magnetic susceptibility, which characterize the LC’s response to external electric and magnetic fields [11]. Mathematically speaking, the LdG $\mathbf{Q}$-tensor is a symmetric, traceless $3\times 3$ matrix, whose eigenvectors model the LC directors, i.e., the special directions of averaged molecular alignment, and whose corresponding eigenvalues are scalar order parameters that quantify the degree of orientational ordering. A LC phase is uniaxial if it has a single distinguished material direction and is biaxial if it has two special directions; more details are given in the next section. The LdG theory has an associated free energy, typically of the form
@@
in the absence of external fields or surface effects. Physically observable configurations are modelled by the LdG energy minimizers subject to the problem constraints. In the case of spatially homogeneous samples, the LdG energy reduces to the bulk potential, $f_{B}$, which is typically a polynomial in the scalar invariants of $\mathbf{Q}$ [17]. The bulk equilibrium phases are then modelled in terms of the minimizers of $f_{B}$, which can be computed explicitly as a function of temperature [23]. The bulk potential, $f_{B}$, is identical for NLC and CLC phases. The elastic energy density, $f_{el}$, depends on both $\mathbf{Q}$ and its gradients, and penalises spatial inhomogeneities. In fact, $f_{el}$, has different forms for the NLC and CLC phases because the natural ground state for a CLC is a helical director profile whereas that for a NLC is a uniform, constant director profile. The critical points and the minimizers of $\pazocal{F}$ are classical solutions to the associated Euler-Lagrange equations: a system of nonlinear, coupled partial differential equations, subject to prescribed constraints.
The question of interest is: how is the LdG energy modified by the NPs in the dilute limit described above? This question has been answered in [7, 8]. In the dilute limit, the NPs do not interact with each other and the NP-LC interaction is captured by a surface energy, $f_{s}$, that depends on the NP geometry and the surface treatment on the NP-LC interface, $\Sigma$. The collective effects of the NPs is, informally speaking, the sum of the surface energies over all NP-LC interfaces which yields a homogenized bulk energy, $f_{hom}(\mathbf{Q})$. More precisely, the total LdG energy, including the effects of the suspended NPs is given by
$\pazocal{F}_{\epsilon}[\mathbf{Q}]:=\pazocal{F}[\mathbf{Q}]+\epsilon^{\gamma}\sum_{i}\int_{\Sigma}f_{s}\left(\mathbf{Q},\mathbf{Q}_{\nu},\boldsymbol{\nu}\right)~dS$
where $\epsilon$ is related to the NP size, $\gamma$ is an appropriate positive scaling factor determined by the dilute limit, $\nu$ is the normal to the NP-LC interface, $\mathbf{Q}_{\nu}$ is a preferred boundary condition on $\Sigma$ and we sum the surface energies over all suspended NPs. In [7], the authors show that minimizers of $\pazocal{F}_{\epsilon}$ converge to minimizers (strongly in $H^{1}$) of
@@
We apply this homogenization framework to a dilute suspension of colloidal NPs in a NLC/CLC sample and analyse the properties of the NP-doped systems in terms of the minimizers of $\pazocal{F}_{0}$, as introduced above. In Section 2, we compute the bulk equilibrium phases of these NP-doped systems as functions of the NP properties and temperature. The bulk phases are modelled by the minimizers of the modified bulk potential, which is the sum of $f_{B}$ and $f_{hom}$. In fact, the bulk equilibrium phases do not distinguish between NLC and CLC samples and hence the conclusions in Section 2 apply to both NLC and CLC samples. The NP properties are captured by two parameters: an anchoring coefficient, $W$, on the NP-LC interfaces and a phenomenological matrix, $\mathbf{X}$, that can be computed from $f_{s}$ directly. We find that NPs stabilise elusive biaxial phases in certain temperature ranges, which are typically inaccessible in pure NLC or CLC samples. Equally importantly, the collective symmetry of the NPs (modelled by $\mathbf{X}$) has a direct bearing on the bulk equilibrium phase at high temperatures and breaks the inherent rotational symmetries of the NLC and CLC phases. These concepts are beautifully illustrated by means of phase diagrams in Section 3, which numerically capture how the bulk equilibrium phases depend on $\mathbf{X}$, $W$ and temperature. In Section 4, we apply this framework to a spatially inhomogeneous CLC-filled channel geometry, doped with spherical NPs in the dilute limit. In the absence of NPs, CLCs naturally relax into twisted helical profiles for relatively large channel widths. The NPs induce untwisting transitions in these geometries as temperature increases, i.e., the NPs naturally break the helical symmetry of the CLC phase. In other words, the NPs collectively impose a special direction (determined by $\mathbf{X}$) on the CLC phase, for sufficiently high temperatures and large channel geometries. These temperature-induced untwisting transitions are specific to CLCs. However, using the same arguments, we do expect NPs to induce symmetry-breaking transitions in other confined LC systems too. We conclude with some perspectives in Section 6.
@@ 2 Homogenized bulk energy
We study a dilute system of suspended colloidal NPs, e.g., quantum dots, in a nematic/cholesteric liquid crystal-filled channel geometry defined by
@@
with $H\ll L\sim D$. We work in the celebrated phenomenological Landau–de Gennes (LdG) framework [10, 11]. The LdG framework describes the macroscopic LC state in terms of the LdG $\mathbf{Q}$-tensor: a symmetric, traceless $3\times 3$ matrix that belongs to the space
@@
We often refer to the eigenvector of $\mathbf{Q}$ with the largest positive eigenvalue as the director, i.e., the molecules tend to align along the director or exhibit long-range orientational ordering along the director, in an averaged sense.
The LC phase is said to be isotropic when $\mathbf{Q}=0$ or equivalently, when the eigenvalues of $\mathbf{Q}$ are zero and there are no special directions/directors. In the uniaxial phase, $\mathbf{Q}$ has two degenerate non-zero eigenvalues and can be written in the form
@@
As with most variational theories in materials science, the nematic/cholesteric phase has an associated LdG free energy and the equilibrium or physically observable configurations are modelled by LdG energy minimizers, subject to the imposed boundary conditions [17]. As mentioned in the Introduction, in the absence of any external fields or surface energies, the LdG energy has two essential contributions:
@@
where $f_{B}$ is the bulk potential, which depends only on the eigenvalues of the $\mathbf{Q}$-tensor and $f_{el}$ is the elastic energy density, which depends on both $\mathbf{Q}$ and its gradient $\nabla\mathbf{Q}$. The bulk potential determines the LC phase (isotropic, uniaxial, or biaxial) in spatially homogeneous samples, as a function of temperature for thermotropic systems. The elastic energy density penalises spatial inhomogeneities stemming from boundary effects, geometric frustration and topological constraints.
@@
where $A=\alpha(T-T^{*})$ is a rescaled temperature, and $B$ and $C$ are positive material-dependent constants. The units of $A,B,C$ are $Nm^{-2}$. Here, $\alpha$ is positive and $T^{*}$ is a critical temperature such that the isotropic phase becomes unstable for $A<0$. As will be shown in the next section, the critical points of $f_{B}$ can be explicitly computed as a function of $T$. The minimizer of $f_{B}$ is the isotropic phase, $\mathbf{Q}=0$, at high temperatures, whereas it is an ordered uniaxial phase for low temperatures. Note that the bulk energy density is identical for nematic and cholesteric liquid crystal phases since cholesterics can be viewed, in non-technical terms, as twisted nematics.
@@
In the homogenization problem, the LC molecules are much smaller than the NPs and the NP-LC interaction can be described by a surface energy. The homogenised energy is the collective contribution of the NP-LC surface energies, summed over all the NPs. More precisely, we model the NP-LC interaction energy per unit area by the familiar weak anchoring energy density [23],
@@
In general, we can treat the $\mathbf{X}$-tensor as a phenomenological $3\times 3$, symmetric, traceless tensor, that is purely determined by the NP-LC surface interactions and represents the collective influence of the spherical NPs, even without knowledge of the underlying microscopic details. For example, if $\mathbf{Q}_{\nu}$ corresponds to homeotropic or radial boundary conditions on a spherical particle, then $\mathbf{X}=0$ [7]. For arbitrary $\mathbf{Q}_{\nu}$ and spherical NPs, $\mathbf{X}$ can be non-trivial and we consider multiple possibilities for $\mathbf{X}$ in the next sections.
@@
where $\pazocal{A}=A/C$, $\pazocal{B}=B/C$ and $W=na^{2}w/C$. As noted in [7, 8], $f_{m}$ differs from $f_{b}$ in two ways - a modified coefficient of $\mathrm{tr}\mathbf{Q}^{2}$ and the addition of a linear forcing term $\mathrm{tr}(\mathbf{Q}\mathbf{X})$. The modified coefficient of $\mathrm{tr}\mathbf{Q}^{2}$ implies that colloidal NPs change the effective temperature of the LC system in the dilute limit. This might be attributed to the effects of the NP-LC interactions which induce some heating or cooling effects depending on the surface treatments on the NP-LC interfaces. One can sometimes view $\mathbf{X}$ as a collective external field and write $\mathbf{X}=\mathbf{E}\otimes\mathbf{E}$. Here, $\mathbf{E}\in\mathbb{R}^{3}$ is an external field and we can redefine $\mathbf{X}=\mathbf{E}\otimes\mathbf{E}-\tfrac{1}{3}\mathbf{I}|\mathbf{E}|^{2}$ to ensure that the isotropic part does not enter into the free energy (8). We use the values $B=3.7\times 10^{5}$ $Nm^{-2}$ and $C=2.4\times 10^{5}$ $Nm^{-2}$ from  [12] and assuming $w=5\times 10^{-3}Nm^{-1}$, $a=50$ nm and $n=10^{22}$$m^{-3}$, we obtain $\pazocal{B}=1.5$ and $W=0.5$, which will be used in the subsequent numerical computations.